Costco has received the following demands for a product in 2020:
Month 1 2 3 4 5 6 7 8 9 10 11 12
Demand 300 700 800 900 3300 200 600 900 200 300 1000 800
Suppose ordering cost (OC) is $504 and holding cost (HC) of one unit of product in a year is $3.
There is no shortage cost. Backordering is not allowed in this model.

Question
Use ‘Lot for Lot’ heuristic method and compute the total cost.

Use ‘Part Period Balancing’ heuristic method, develop a schedule to show when and how many products should be
ordered, and compute the total cost.
Note: to compute holding cost, use average inventory (not ending inventory).

Use ‘Silver Meal’ heuristic method, develop a schedule to show when and how many products should be ordered, and
compute the total cost.
Note: to compute holding cost, use average inventory (not ending inventory).
Answer #1

1) Lot for Lot

Using 'Lot for Lot' heuristic, order quantity is equal to demand in each period. So, starting inventory for each month is equal to monthly demand and ending inventory for each month is equal to 0. Average inventory = (starting inventory + ending inventory)/2 = Monthly demand / 2

Sum of average inventory of 12 months = (300+700+800+900+3300+200+600+900+200+300+1000+800)/2 = 5000

Number of orders = 12

Total cost = 12*OC + (5000/12)*HC = 12*504+(5000/12)*3 = $ 7,298

2) Part Period Balancing

In this method, a basic unit of PP (Part-Period) is used. When one part is stored for one period, the value is 1 PP, and the total value of all PP values calculated for each period is an accumulated PP. When accumulated inventory carrying cost excesses the order cost, the quantity of schedule order is set. In order to make these calculations easier, a standard value for lot sizing, called EPP (Economic Part-Period), is used. This value, meaning that the accumulated PP is over the order cost, is calculated by means of the following formula:
EPP Value = Setup cost / Inventory holding cost

Holding cost per month = 3/12 = $ 0.25

3) Silver Meal heuristic

The sum of ordering and holding cost for multiple trial lot sizes is computed and that total cost is divided by the number of periods that the trial lot size covers to get the cost per period C(t). The trial lot size, for which this cost per period, C(t) is the lowest is selected. This process is repeated for future demands, which are not yet covered by the previous trial lot size.